Please help me solving this word problem. The question is below---
A card collecting store sells individual packs for one price or an entire hobby box for another. Jaycee purchased 12 packs and 3 hobby boxes for $196.50. Jeremiah paid $403.50 for 8 packs and 7 hobby boxes.

Answer by stanbon(75887)

A card collecting store sells individual packs for one price or an entire hobby box for another. Jaycee purchased 12 packs and 3 hobby boxes for $196.50. Jeremiah paid $403.50 for 8 packs and 7 hobby boxes.
----------
Equations:
p is the price of a pack ; h is the price of a hobby box
-------
Value Eq: 12p + 3h = 196.50
Value Eq: 8p + 7h = 403.50
------------------------------------
Modify:
96p + 24h = 8*196.50
96p + 84h = 12*403.50
---------------------------
60h = 3270
h = $54.50 (price of a hobby box)
------
Solve for "p":
8p + 7h = 403.50
8p + 381.50 = 403.50
8p = 22
p = $2.75 (price of a pack)
=================================
Cheers,
Stan H.
========================